OpenStudy (anonymous):

The domain of the given function is {4, -2}. What is the range?

OpenStudy (anonymous):

h (x) = 2x^2 - 5

Parth (parthkohli):

You plug-in 4 and -2 in the function respectively.

Parth (parthkohli):

Essentially, you have to find \(h(4)\) and \(h(-2)\).

Parth (parthkohli):

Can you do that?

OpenStudy (anonymous):

ooo ok let me try

Parth (parthkohli):

Sure

OpenStudy (anonymous):

h(4) = 2(4) and does h (-2) = 2(-2)

Parth (parthkohli):

No...

Parth (parthkohli):

\[ h(4) = 2(4)^2 - 5\\h(-2) = 2(-2)^2 - 5\]

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

so now i have to multiply these or solve i mean

OpenStudy (anonymous):

is the final answer (-13, 11)

OpenStudy (anonymous):

hello?

Parth (parthkohli):

Nope, not.

OpenStudy (anonymous):

how come?

OpenStudy (anonymous):

i did the work and thats what i found out though

OpenStudy (anonymous):

or maybe i did it wrong

Parth (parthkohli):

\[h(4) =2(4)^2 - 5 \implies 2(16) - 5 \implies 32 - 5 = 27 \]

OpenStudy (anonymous):

ooooooooooo ok

OpenStudy (anonymous):

i got my answer now, is it 3,27

Parth (parthkohli):

No!

Parth (parthkohli):

You got the 3 and 27 part right, but since the domain is {4,-2}, we have {27,3} as the range.

OpenStudy (anonymous):

o ok i get it thnx

Parth (parthkohli):

You're welcome! And were you born in 1999 too?

OpenStudy (anonymous):

yes...

Parth (parthkohli):

Me too :)

OpenStudy (anonymous):

nice!

OpenStudy (anonymous):

can u help me with one more question?

OpenStudy (anonymous):

18 - 8x - 15 = -7 - 5 ( x - 1) + 5

OpenStudy (anonymous):

i got the answer -5x = 7

Parth (parthkohli):

Can you post it after closing this one?

OpenStudy (anonymous):

sure